# Graph Optimality-Aware Stochastic LiDAR Bundle Adjustment with Progressive Spatial Smoothing

### v1

@@ Abstract
Large-scale LiDAR Bundle Adjustment (LBA) for refining sensor orientation and point cloud accuracy simultaneously is a fundamental task in photogrammetry and robotics, particularly as low-cost 3D sensors are increasingly used for 3D mapping in complex scenes. Unlike pose-graph-based methods that rely solely on pairwise relationships between LiDAR frames, LBA leverages raw LiDAR correspondences to achieve more precise results, especially when initial pose estimates are unreliable for low-cost sensors. However, existing LBA methods face challenges such as simplistic planar correspondences, extensive observations, and dense normal matrices in the least-squares problem, which limit robustness, efficiency, and scalability. To address these issues, we propose a Graph Optimality-aware Stochastic Optimization scheme with Progressive Spatial Smoothing, namely PSS-GOSO, to achieve robust, efficient, and scalable LBA. The Progressive Spatial Smoothing (PSS) module extracts robust LiDAR feature association exploiting the prior structure information obtained by the polynomial smooth kernel. The Graph Optimality-aware Stochastic Optimization (GOSO) module first sparsifies the graph according to optimality for an efficient optimization. GOSO then utilizes stochastic clustering and graph marginalization to solve the large-scale state estimation problem for a scalable LBA. We validate PSS-GOSO across diverse scenes captured by various platforms, demonstrating its superior performance compared to existing methods.
keywords:
@@
LiDAR Bundle Adjustment (LBA) , Point Clouds , Stochastic Optimization , 3D Mapping
journal: Journal of Photogrammetry and Remote Sensing
1 Introduction
@@
LiDAR Bundle Adjustment (LBA) is a fundamental task in photogrammetry (Habib et al., 2009) and robotics (Liu et al., 2023), aimed at simultaneously refining sensor orientation and point cloud accuracy. As a critical preprocessing step, LBA reduces point cloud discrepancies and enhances consistency, which is essential for downstream applications such as forest inventory (Li et al., 2019), infrastructure maintenance (Lin and Habib, 2022), and robotic delivery utilizing prior maps (Li et al., 2024c).
@@
Figure 1: Graph optimality-aware stochastic optimization with progressive spatial smoothing for a robust, efficient, and scalable LBA. The Progressive Spatial Smoothing (PSS) module extracts robust LiDAR feature association exploiting the prior structure information obtained by the polynomial smooth kernel. The Graph Optimality-aware Stochastic Optimization (GOSO) module first sparisifies the graph according to optimality for an efficient optimization. GOSO then utilizes stochastic clustering and graph marginalization to solve the large-scale state estimation problem for a scalable LBA.
@@
LBA has been extensively studied over the past decades and has been applied in traditional mobile mapping systems, such as aerial laser scanning (Glira et al., 2015) and vehicle-borne laser scanning (Li et al., 2023b). However, with the rise of compact, low-cost 3D sensing systems and the increasing demand for 3D mapping in GNSS-denied, complex environments, existing LBA methods are not directly applicable to these low-cost platforms due to issues related to robustness, efficiency, and computational load. Consequently, adapting LBA for compact sensing systems in large-scale, complex scenes has become a prominent topic of interest in both academic and industrial communities.
@@
1.1 Indirect optimization-based method
@@
Indirect optimization-based methods first extract relative transformations by conducting pair-wise registration techniques, such as ICP (Segal et al., 2009), NDT (Einhorn and Gross, 2015), and their variants (Koide et al., 2021), and then optimize the poses without involving the raw correspondences used in the pair-wise registration. For example, Dong et al. (2018) registers unordered terrestrial laser scanning point clouds by analyzing the pair-wise relationships between laser frames, integrating the point clouds incrementally without redundant pose constraints. This approach can be seen as one of the simplest indirect optimization-based methods for fusing multiple point clouds. However, the incremental fusion approach may lead to error accumulation, which is often mitigated using pose graph optimization (Carlone et al., 2014). A critical challenge in accurate pose graph optimization is transforming the pair-wise registration results into relative pose constraints that follow a Gaussian distribution (Yang and Li, 2022). In practice, this Gaussian assumption is often overestimated (Koide et al., 2022), which can result in information loss and inaccuracies. Additionally, pose graph optimization may produce erroneous results if the relative pose estimation is unreliable in areas with low frame overlap. The indirect optimization-based methods are not the primary focus of this paper.
1.2 Direct optimization-based method
@@
Direct optimization-based methods extract feature correspondences between LiDAR frames and simultaneously optimize both the poses and the point clouds. Compared to indirect optimization-based methods, direct optimization often achieves higher accuracy by avoiding information loss during the relative transformation estimation step inherent in indirect methods. However, three main challenges need to be addressed in direct optimization-based methods: feature association, efficiency, and scalability.
@@
Feature association in LBA determines the type of correspondences used for pose correction. Commonly utilized features include points (Li et al., 2023b), lines (Lee et al., 2007), planes (Skaloud and Lichti, 2006), and their combinations (Liu and Zhang, 2021). While these features perform effectively in most structured environments, their performance can degrade and lead to divergence in complex environments that lack clear structural features. Despite the effectiveness of these commonly used features, our previous work, PSS-BA (Li et al., 2024b), demonstrated that spatial smoothing can enhance feature association by providing more robust and richer constraints in complex scenarios. Nevertheless, PSS-BA (Li et al., 2024b) is constrained by data volume limitations and is not well-suited for large-scale environments. Consequently, this work aims to extend PSS-BA to enhance its efficiency and scalability.
@@
Efficiency is crucial for end users, particularly in large-scale mapping tasks. Many existing pose optimization techniques face a trade-off between accuracy and efficiency (Li et al., 2024c; Jiao et al., 2021). For instance, Jiang and Jiang (2017) accelerates image bundle adjustment by leveraging the graph structure of poses and employing maximal spanning tree expansion. Similarly, Li et al. (2024c) enhances real-time performance in continuous state estimation by optimizing the normal matrix. However, to the best of our knowledge, no current LBA methods effectively select the most useful constraints from redundant correspondences to improve the efficiency of the optimization process while maintaining comparable accuracy.
@@
Scalability is a critical issue for LiDAR Bundle Adjustment (LBA), especially in large-scale mapping applications. Traditional mobile mapping systems equipped with high-accuracy Positioning and Orientation Systems (POS) that work with GNSS signals can model the small orientation errors from the high-accuracy POS as splines for correction (Pöppl et al., 2024). In such cases, a relatively small number of feature correspondences can suffice for achieving high-quality results in large-scale applications (Brun et al., 2022). The sparse normal matrix used in the least-square estimation resulting from a small number of features is manageable with sparse-solving tools (Davis, 2019). However, with the rise of low-cost sensors used for 3D mapping in GNSS-denied complex environments, scalability becomes more challenging. Low-cost LiDAR SLAM often provides less accurate initial poses, whose errors cannot be effectively modeled using splines to reduce system states. Additionally, the dense correspondences used for optimization result in a very dense normal matrix, which limits the scalability of existing LBA methods. To address these issues, Yang and Li (2022) introduced a scan-block structure to segment LiDAR frames from UAV systems, improving data quality. Similarly, Liu et al. (2023) proposed a hierarchical adjustment scheme that splits LiDAR frames into blocks based on time, thus mitigating data volume limitations. However, these sequential splitting strategies (Yang and Li, 2022; Liu et al., 2023) do not fully account for the spatial relationships in the pose graph and suffer from non-rigid drift in local blocks, leading to discrepancies in the final point clouds (Li et al., 2023a). Moreover, such sequential splitting methods are not well-suited for multi-session tasks, such as integrating scanning data collected from different platforms for large-scale scenes (Nair et al., 2024).
@@
The remainder of this paper is organized as follows: Section 2 provides the preliminary details for our system. The Progressive Spatial Smoothing (PSS) module, which is used for robust LiDAR feature association, is described in Section 4. Section 5 presents a detailed description of the Graph Optimality-aware Stochastic Optimization (GOSO) method. Section 6 covers the experiments conducted using both public and proprietary datasets. Finally, Section 7 offers conclusions and discusses future work.
2 Preliminary
2.1 Notation
2.1.1 Poses
In this paper, we use italic, bold lowercase, and bold uppercase letters to represent scalars, vectors, and matrices, respectively. Three main frames of reference are used in our proposed method, namely, the world frame $F^{W}$, the LiDAR frame $F^{L}$, and the tangent frame $F^{S}$ defined in a local 3D space (Dong et al., 2017). We denote any point observed by the LiDAR in the $i$-th frame as $\mathbf{p}^{L}_{i}$. The pose for the $i$-th LiDAR frame is denoted as $\mathbf{X}_{i}=[\mathbf{R}_{i},\mathbf{t}_{i}]$, where $\mathbf{R}_{i}\in\mathrm{SO(3)}$ is the rotation matrix, $\mathbf{t}_{i}\in\mathbb{R}^{3}$ is the translation vector. All the poses are denoted as $\mathbf{X}=\{\mathbf{X}_{i}\}^{m}_{i=1}$. The initial value is noted with breve $\breve{\circ}$, and the estimated value is noted with the hat $\hat{\circ}$. The estimated pose of the $i$-th LiDAR frame is denoted as $\hat{\mathbf{X}}_{i}=[\hat{\mathbf{R}}_{i}=\mathrm{Exp}{(\Delta\bm{\theta}_{i})}\mathbf{R},\hat{\mathbf{t}}_{i}=\Delta\mathbf{t}_{i}+\mathbf{t}_{i}]$, where $\Delta\bm{\theta}_{i}\in\mathbb{R}^{3}$ and $\Delta\mathbf{t}_{i}\in\mathbb{R}^{3}$ are the corresponding errors. The exponential map $\mathrm{Exp}:\mathbb{R}^{3}\rightarrow\mathrm{SO(3)}$ has the form:
@@ 2.1.1 Poses
2.1.3 Polynomial Smoothing Kernel
@@
2.2 LiDAR Bundle Adjustment
A LBA problem can be built upon a relation graph $\mathbf{G}=(\mathbf{X},\mathbf{E})$. Here, $\mathbf{X}$ is the set of $m$ LiDAR poses. $\mathbf{E}=\{\mathbf{E}_{k}\}_{k=1}^{l}$ denotes the $l$ correspondences between the LiDAR poses. The feature correspondences can be the point-to-point, point-to-line, and point-to-planarity styles (Zhang and Singh, 2014; Segal et al., 2009). In this work, we use the polynomial smoothing kernel (4) to formulate the correspondences, which will be detailed in Section 4. Generally, the feature correspondence $\mathbf{E}_{k}$ depends on the actual poses $\mathbf{X}$, i.e. $\mathbf{E}_{k}=\mathrm{E}(\mathbf{X})_{k}$. In essence, LBA is the following minimization problem (5).
@@ 2.2 LiDAR Bundle Adjustment
where $\mathrm{E}(\mathbf{X})$ is the concatnation of $\{\mathrm{E}(\mathbf{X})_{k}\}^{l}_{k=1}$. Setting the value of $\mathbf{X}$ with the initial guess $\mathbf{\mathbf{\breve{X}}}$, the Levenberg Marquardt (LM) algorithm (Ranganathan, 2004) achieves a pose correction step $\Delta\mathbf{X}$ at each iteration by linearizing $\mathrm{F}(\mathbf{\mathbf{\breve{X}}})$ as (6).
$\underset{\Delta\mathbf{X}}{\mathrm{argmin}}\leavevmode\nobreak\ \mathrm{F}(\mathbf{X})=\|\mathbf{J}\Delta\mathbf{X}+\mathrm{E}(\mathbf{\breve{X}})\|^{2}+\lambda\|\mathbf{D}\Delta\mathbf{X}\|^{2},$ (6)
@@
2.3 Optimality of LiDAR Bundle Adjustment
Without solving the full bundle adjustment problem (7), we could still obtain the expected accuracy of (7) by analyzing the optimality of the relation graph $\mathbf{G}$. For a random edge $\mathbf{E}_{k}$, first and second elements are the $\mathbf{E}_{k}^{0}$-th pose and $\mathbf{E}_{k}^{1}$-th pose, respectively. The relative pose residual ($\bm{\epsilon}_{k}$) and covariance ($\mathbf{\Omega}_{k}$) between the $\mathbf{E}_{k}^{0}$-th pose and $\mathbf{E}_{k}^{1}$-th pose are approximated using the pair-wise registration formulation in B. The minimal eigen value of the covariance $\lambda^{min}_{k}=\lambda_{min}(\bm{\mathbf{\Omega}}_{k})$ is used to measure the reliability of the edge $\mathbf{E}_{k}$. The optimality of $\mathbf{G}$ is first approximated by formulating the standard pose graph equation (8),
$\underset{\mathbf{X}}{\mathrm{argmin}}\leavevmode\nobreak\ \sum_{k=1}^{l}\bm{\epsilon}_{k}^{\top}\mathbf{\Omega}_{k}\bm{\epsilon}_{k}.$ (8)

### v2

@@ Abstract
Large-scale LiDAR Bundle Adjustment (LBA) to refine sensor orientation and point cloud accuracy simultaneously to build the navigation map is a fundamental task in logistics and robotics. Unlike pose-graph-based methods that rely solely on pairwise relationships between LiDAR frames, LBA leverages raw LiDAR correspondences to achieve more precise results, especially when initial pose estimates are unreliable for low-cost sensors. However, existing LBA methods face challenges such as simplistic planar correspondences, extensive observations, and dense normal matrices in the least-squares problem, which limit robustness, efficiency, and scalability. To address these issues, we propose a Graph Optimality-aware Stochastic Optimization scheme with Progressive Spatial Smoothing, namely PSS-GOSO, to achieve robust, efficient, and scalable LBA. The Progressive Spatial Smoothing (PSS) module extracts robust LiDAR feature association exploiting the prior structure information obtained by the polynomial smooth kernel. The Graph Optimality-aware Stochastic Optimization (GOSO) module first sparsifies the graph according to optimality for an efficient optimization. GOSO then utilizes stochastic clustering and graph marginalization to solve the large-scale state estimation problem for a scalable LBA. We validate PSS-GOSO across diverse scenes captured by various platforms, demonstrating its superior performance compared to existing methods. Moreover, the resulting point cloud maps are used for automatic last-mile delivery in large-scale complex scenes. The project page can be found at: https://kafeiyin00.github.io/PSS-GOSO/.
Index Terms:
@@
LiDAR Bundle Adjustment (LBA); Point Clouds; Stochastic Optimization; 3D Mapping
I Introduction
@@
LiDAR Bundle Adjustment (LBA) is a fundamental task in logistic [1] and robotics [2], aiming to simultaneously refine sensor orientation and point cloud map accuracy. As a critical step for accurate 3D mapping of the complex scenes, LBA reduces point cloud discrepancies and enhances consistency, which is essential for downstream applications such as infrastructure maintenance [3, 4], and robotic last-mile delivery utilizing prior maps [5, 6].
@@
Figure 1: Graph optimality-aware stochastic optimization with progressive spatial smoothing for a robust, efficient, and scalable LBA. The Progressive Spatial Smoothing (PSS) module extracts robust LiDAR feature association exploiting the prior structure information obtained by the polynomial smooth kernel. The Graph Optimality-aware Stochastic Optimization (GOSO) module first satisfies the graph according to optimality for an efficient optimization. GOSO then utilizes stochastic clustering and graph marginalization to solve the large-scale state estimation problem for a scalable LBA.
@@
LBA has been extensively studied over the past decades and has been applied in traditional mobile mapping systems, such as aerial laser scanning [7] and vehicle-borne laser scanning [8]. However, with the rise of compact, low-cost 3D sensing systems and the increasing demand for 3D mapping in complex environments denied by GNSS, existing LBA methods are not directly applicable to these low-cost platforms due to issues related to robustness, efficiency, and computational load. Consequently, adapting LBA for compact sensing systems in large-scale, complex scenes has become a prominent topic of interest in both academic and industrial communities.
@@
I-A Indirect optimization-based method
@@
Indirect optimization-based methods first extract relative transformations by performing pairwise registration techniques, such as ICP [9], NDT [10], and their variants [11], and then optimize poses without involving the raw correspondences used in pairwise registration. For example, [12] registers unordered terrestrial laser scanning point clouds by analyzing the pair-wise relationships between laser frames, integrating the point clouds incrementally without redundant pose constraints. This approach can be seen as one of the simplest indirect optimization-based methods to merge multiple point clouds. However, the incremental fusion approach may lead to error accumulation, which is often mitigated by using pose graph optimization [13]. A critical challenge in accurate pose graph optimization is transforming pairwise registration results into relative pose constraints that follow a Gaussian distribution [14]. In practice, this Gaussian assumption is often overestimated [15], which can result in information loss and inaccuracies. Additionally, pose graph optimization may produce erroneous results if the relative pose estimation is unreliable in areas with low frame overlap. Indirect optimization-based methods are not the primary focus of this paper.
I-B Direct optimization-based method
@@
Direct optimization-based methods extract feature correspondences between LiDAR frames and simultaneously optimize both poses and point clouds. Compared to indirect optimization-based methods, direct optimization often achieves higher accuracy by avoiding information loss during the relative transformation estimation step inherent in indirect methods. However, three main challenges need to be addressed in direct optimization-based methods: feature association, efficiency, and scalability.
@@
Feature association in LBA determines the type of correspondence used for pose correction. Commonly utilized features include points [8], planes [16], and their combinations [17]. While these features perform effectively in most structured environments, their performance can degrade and lead to divergence in complex environments that lack clear structural features. Despite the effectiveness of these commonly used features, existing work, PSS-BA [18], demonstrated that spatial smoothing can enhance feature association by providing more robust and richer constraints in complex scenarios. Nevertheless, PSS-BA [18] is constrained by data volume limitations and is not well-suited for large-scale environments. Consequently, this work aims to extend PSS-BA to enhance its efficiency and scalability.
@@
Efficiency is crucial for end users, particularly in large-scale mapping tasks. Many existing pose optimization techniques face a trade-off between accuracy and efficiency [6]. For example, [19] accelerates the adjustment of the image bundle by taking advantage of the graph structure of poses and employing the maximal expansion of the spanning tree. Similarly, [6] enhances real-time performance in continuous state estimation by optimizing the normal matrix. However, to the best of our knowledge, no current LBA methods effectively select the most useful constraints from redundant correspondences to improve the efficiency of the optimization process while maintaining comparable accuracy.
@@
Scalability is a critical issue for LiDAR Bundle Adjustment (LBA), especially in large-scale mapping applications. Traditional mobile mapping systems equipped with high-accuracy Positioning and Orientation Systems (POS) that work with GNSS signals can model the small orientation errors from the high-accuracy POS as splines for correction [20]. In such cases, a relatively small number of feature correspondences can suffice for achieving high-quality results in large-scale applications. The sparse normal matrix used in the least-square estimation resulting from a small number of features is manageable with sparse-solving tools. However, with the rise of low-cost sensors used for 3D mapping in complex environments denied GNSS, scalability becomes more challenging. Low-cost LiDAR SLAM often provides less accurate initial poses, whose errors cannot be effectively modeled using splines to reduce system states. Additionally, the dense correspondences used for optimization result in a very dense normal matrix, which limits the scalability of existing LBA methods. To address these issues, [14] introduced a scan-block structure to segment LiDAR frames from UAV systems, improving data quality. Similarly, [2] proposed a hierarchical adjustment scheme that splits LiDAR frames into blocks based on time, thus mitigating data volume limitations. However, these sequential splitting strategies [2] do not fully account for the spatial relationships in the pose graph and suffer from non-rigid drift in local blocks, leading to discrepancies in the final point clouds [21]. Moreover, such sequential splitting methods are not well-suited for multi-session tasks, such as integrating scanning data collected from different platforms for large-scale scenes [22].
@@
The remainder of this paper is organized as follows: Section II provides the preliminary details for our system. The Progressive Spatial Smoothing (PSS) module, which is used for robust LiDAR feature association, is described in Section IV. Section V presents a detailed description of the Graph Optimality-aware Stochastic Optimization (GOSO) method. Section VI covers the experiments conducted using both public and proprietary datasets. Finally, Section VII offers conclusions and discusses future work.
II Preliminary
II-A Notation
II-A1 Poses
In this paper, we use italic, bold lowercase, and bold uppercase letters to represent scalars, vectors, and matrices, respectively. Three main frames of reference are used in our proposed method, namely, the world frame $F^{W}$, the LiDAR frame $F^{L}$, and the tangent frame $F^{S}$ defined in a local 3D space [23]. We denote any point observed by the LiDAR in the $i$-th frame as $\mathbf{p}^{L}_{i}$. The pose for the $i$-th LiDAR frame is denoted as $\mathbf{X}_{i}=[\mathbf{R}_{i},\mathbf{t}_{i}]$, where $\mathbf{R}_{i}\in\mathrm{SO(3)}$ is the rotation matrix, $\mathbf{t}_{i}\in\mathbb{R}^{3}$ is the translation vector. All the poses are denoted as $\mathbf{X}=\{\mathbf{X}_{i}\}^{m}_{i=1}$. The initial value is noted with breve $\breve{\circ}$, and the estimated value is noted with the hat $\hat{\circ}$. The estimated pose of the $i$-th LiDAR frame is denoted as $\hat{\mathbf{X}}_{i}=[\hat{\mathbf{R}}_{i}=\mathrm{Exp}{(\Delta\bm{\theta}_{i})}\mathbf{R},\hat{\mathbf{t}}_{i}=\Delta\mathbf{t}_{i}+\mathbf{t}_{i}]$, where $\Delta\bm{\theta}_{i}\in\mathbb{R}^{3}$ and $\Delta\mathbf{t}_{i}\in\mathbb{R}^{3}$ are the corresponding errors. The exponential map $\mathrm{Exp}:\mathbb{R}^{3}\rightarrow\mathrm{SO(3)}$ has the form:
@@ II-A1 Poses
II-A3 Polynomial Smoothing Kernel
@@
II-B LiDAR Bundle Adjustment
A LBA problem can be built upon a relation graph $\mathbf{G}=(\mathbf{X},\mathbf{E})$. Here, $\mathbf{X}$ is the set of $m$ LiDAR poses. $\mathbf{E}=\{\mathbf{E}_{k}\}_{k=1}^{l}$ denotes the $l$ correspondences between the LiDAR poses. The feature correspondences can be the point-to-point, point-to-line, and point-to-planarity styles [24, 9]. In this work, we use the polynomial smoothing kernel (4) to formulate the correspondences, which will be detailed in Section IV. Generally, the feature correspondence $\mathbf{E}_{k}$ depends on the actual poses $\mathbf{X}$, i.e. $\mathbf{E}_{k}=\mathrm{E}(\mathbf{X})_{k}$. In essence, LBA is the following minimization problem (5).
@@ II-B LiDAR Bundle Adjustment
where $\mathrm{E}(\mathbf{X})$ is the concatnation of $\{\mathrm{E}(\mathbf{X})_{k}\}^{l}_{k=1}$. Setting the value of $\mathbf{X}$ with the initial guess $\mathbf{\mathbf{\breve{X}}}$, the Levenberg Marquardt (LM) algorithm [25] achieves a pose correction step $\Delta\mathbf{X}$ at each iteration by linearizing $\mathrm{F}(\mathbf{\mathbf{\breve{X}}})$ as (6).
$\underset{\Delta\mathbf{X}}{\mathrm{argmin}}~{}\mathrm{F}(\mathbf{X})=\|\mathbf{J}\Delta\mathbf{X}+\mathrm{E}(\mathbf{\breve{X}})\|^{2}+\lambda\|\mathbf{D}\Delta\mathbf{X}\|^{2},$ (6)
@@
II-C Optimality of LiDAR Bundle Adjustment
Without solving the full bundle adjustment problem (7), we could still obtain the expected accuracy of (7) by analyzing the optimality of the relation graph $\mathbf{G}$. For a random edge $\mathbf{E}_{k}$, first and second elements are the $\mathbf{E}_{k}^{0}$-th pose and $\mathbf{E}_{k}^{1}$-th pose, respectively. The relative pose residual ($\bm{\epsilon}_{k}$) and covariance ($\mathbf{\Omega}_{k}$) between the $\mathbf{E}_{k}^{0}$-th pose and $\mathbf{E}_{k}^{1}$-th pose are approximated using the pair-wise registration formulation in Appendix B: Relative Pose Residual and Covariance from Pair-wise Point Cloud Registration. The minimal eigen value of the covariance $\lambda^{min}_{k}=\lambda_{min}(\bm{\mathbf{\Omega}}_{k})$ is used to measure the reliability of the edge $\mathbf{E}_{k}$. The optimality of $\mathbf{G}$ is first approximated by formulating the standard pose graph equation (8),
$\underset{\mathbf{X}}{\mathrm{argmin}}~{}\sum_{k=1}^{l}\bm{\epsilon}_{k}^{\top}\mathbf{\Omega}_{k}\bm{\epsilon}_{k}.$ (8)